OpenStudy (anonymous):

Solve 5c - c + 10 = 34. (1 point) -6 6 11 -11

OpenStudy (anonymous):

6

OpenStudy (anonymous):

4c=24=> c=6

OpenStudy (nurali):

5c - c + 10 = 34 4c+10=34 4c=34-10 4c=24 c=24/4 c=6

OpenStudy (anonymous):

Thank you both kindly :)

OpenStudy (anonymous):